Application, I am having a hard time setting this problem up.
The snack shop makes three mixes of nuts in the following proportions:
Mix I: 7 pounds peanuts, 5 pounds cashews, and 1 pound pecans.
Mix II: 3 pounds peanuts, 2 pounds cashews and 2 pounds pecans.
Mix III: 4 pounds peanuts, 3 pounds cashews, and 3 punds pecans.
The shop has 67 pounds of peanuts, 48 pounds of cashews and 32 pounds of pecans in stock. How many of each mix can it produce?

Answer by ewatrrr(10682)

 
Hi,
Let x,y,z be the number of each kind of mix that can be made
that is:
7x+3y+4z = 67
5x+2y+3z = 48
x+2y+3z = 32
solve the 3 Eqs in the manner you choose
x = 4, y = 5, z = 6